A random sample of n=25 individuals is selected from a population with mean =20, and a treatment is administered to each individual sample. After treatment, the sample mean is found to be M=22.2 with as=384. If there is no treatment effect, how much difference is expected between the sample mean and it's population mean. Find the standard error for M.
Answer #1

Standard error of M

u/s =

= 384/ 25

= 384/5

= 76.8